OpenStudy (anonymous):

Gina is deciding between two new cars. The non-hybrid model she likes costs $14,000 plus $0.23 per mile. The hybrid model she likes costs $24,000 plus $0.13 per mile. She plans to buy another new car when the odometer reaches 100,000 miles. Graph the system of linear equations to determine which car is a better financial choice. The hybrid vehicle is a better value. The standard vehicle is a better value. Either − after that amount of time, she will have spent the same amount. Not enough information.

OpenStudy (anonymous):

Write out an equation for both scenarios. 4th sentence is irrelevant information. Total cost of non-hybrid = 14,000 + 0.23x x is number of miles. Total cost of hybrid = 24,000 + 0.13x Not enough information, because they don't tell you how many miles she plans to drive in the NEW car. They tell you how long she's going to wait to buy a new car but that's it.

OpenStudy (anonymous):

Write out an equation for both scenarios. 4th sentence is irrelevant information. Total cost of non-hybrid = 14,000 + 0.23x x is number of miles. Total cost of hybrid = 24,000 + 0.13x Not enough information, because they don't tell you how many miles she plans to drive in the NEW car. They tell you how long she's going to wait to buy a new car but that's it.